# Zak-OTFS with Interleaved Pilots to Extend the Region of Predictable Operation

### v1

@@ VI Numerical simulations
We report simulation results for the Veh-A channel model [22] which consists of six channel paths. The channel gains $h_{i}$ are modeled as independent zero-mean complex circularly symmetric Gaussian random variables, normalized so that $\sum\limits_{i=1}^{6}{\mathbb{E}}\left[|h_{i}|^{2}\right]=1$. Table I lists the power-delay profile for the six channel paths. The Doppler shift of the $i$-th path is modeled as $\nu_{i}=\nu_{max}\,\cos\left(\theta_{i}\right)$, where $\nu_{max}$ is the maximum Doppler shift of any path, and the variables $\theta_{i},i=1,2,\cdots,6$, are independent and uniformly distributed in $[-\pi\,,\,\pi)$.
@@
where $0\leq\beta_{\tau},\beta_{\nu}\leq 1$ and
@@
We employ the matched filter $w_{rx}(\tau,\nu)=w_{tx}^{*}(-\tau,-\nu)\,e^{j2\pi\nu\tau}$ at the receiver. We choose roll-off factors $\beta_{\nu}=\beta_{\tau}=0.6$ so that the effective bandwidth of the Zak-OTFS subframe is $B^{\prime}=(1+\beta_{\tau})B$ and the effective time duration is $T^{\prime}=(1+\beta_{\nu})T$.
@@
In Fig. 12 we plot the effective throughput as a function of increasing $\nu_{max}$ for the same simulation setting as in Fig. 10. Effective throughput is the ratio of the number of bits reliably communicated in each subframe to the available degrees of freedom ($B^{\prime}T^{\prime}$). The number of bits communicated reliably in a subframe is simply $(1-H(BER))$ times the number of information bits transmitted in each subframe. Here BER denotes the bit error rate and $H(\cdot)$ denotes the binary entropy function. We maximize effective throughput by interleaving the minimum number of pilots required to accurately estimate the effective channel. We minimize the number of pilots to avoid introducing unnecessary guard and pilot regions that would reduce effective throughput. When $2\nu_{max}<\nu_{p}$ we use a single pilot, when $\nu_{p}<2\nu_{max}<2\nu_{p}$ we use $2$ interleaved pilots, and when $2\nu_{p}<2\nu_{max}<4\nu_{p}$ we use $4$ interleaved pilots. Note that although a higher number of interleaved pilots results in stable throughput for a wider range of Doppler spreads (i.e., extension of the region of predictable operation), the throughput achieved is smaller due to a higher pilot and guard region overhead.
Figure 13: BER vs PDR for six-path Veh-A channel used in Fig. 10 and 12, with a fixed $\nu_{max}=2.5$ KHz.
@@
The peak to average power ratio (PAPR) of the transmitted TD pilot (no data transmission) depends on the number of interleaved pilots. For the Zak-OTFS system considered here, the PAPR decreases from $19.4$ dB, to $16.4$ dB to $13.4$ dB as the number of interleaved pilots increases from $1$, to $2$, to $4$. This reduction is illustrated in Fig. 14. When the number of interleaved pilots doubles, the separation between pulses in the TD pilot pulse train halves. [^11] The TD realization of a single impulse pilot at DD location $(k_{p_{1}},0)$ is a TD pulse train with narrow TD pulses at time instances $\left(\frac{k_{p_{1}}\tau_{p}}{M}+n\tau_{p}\right)$, $n=-N/2,\cdots,0,(N/2-1)$ for even $N$ (see [18] and [1]). Therefore, the TD realization of $Q$ interleaved pilots spaced regularly apart at DD locations $(iM/Q,0)$, $i=0,1,\cdots,(Q-1)$ is a superposition of $Q$ pulse trains, each pulse train being the TD realization to one of the $Q$ pilots. The TD realization of $Q$ interleaved pilots is therefore a pulse train consisting of narrow TD pulses spaced $\tau_{p}/Q$ seconds apart. For the same total pilot energy $E_{p}$, the energy of each narrow TD pulse is therefore $E_{p}/Q$. There are twice as many pulses, and each pulse is scaled down by $\sqrt{2}$ to maintain constant average power. Hence the PAPR is halved with every doubling of the number of interleaved pilots.

### v2

@@ VI Numerical simulations
We report simulation results for the Veh-A channel model [25] which consists of six channel paths. The channel gains $h_{i}$ are modeled as independent zero-mean complex circularly symmetric Gaussian random variables, normalized so that $\sum\limits_{i=1}^{6}{\mathbb{E}}\left[|h_{i}|^{2}\right]=1$. Table I lists the power-delay profile for the six channel paths. The Doppler shift of the $i$-th path is modeled as $\nu_{i}=\nu_{max}\,\cos\left(\theta_{i}\right)$, where $\nu_{max}$ is the maximum Doppler shift of any path, and the variables $\theta_{i},i=1,2,\cdots,6$, are independent and uniformly distributed in $[-\pi\,,\,\pi)$. [^13] We consider fractional delay and Doppler shifts, which is representative of real propagation environments. Note that path delays in Table-I are non-integer multiples of the delay domain resolution $1/B$. The Doppler shifts $\nu_{i}=\nu_{max}\cos(\theta_{i})$ are also non-integer multiples of the Doppler domain resolution $1/T$ since $\cos(\theta_{i})$ is continuous valued. Also, the proposed scheme of using interleaved pilots for acquisition of the I/O relation is still applicable if we have two paths with same delay but different Doppler shifts, since the required number of interleaved pilots depends only on the maximum spread of $h_{\mbox{\scriptsize{eff}}}[k,l]$ along the Doppler axis.
@@
where $0\leq\beta_{\tau},\beta_{\nu}\leq 1$ and $rrc_{{}_{\beta}}(x)$ is the RRC waveform (see [23]) given by
@@
We employ the matched filter $w_{rx}(\tau,\nu)=w_{tx}^{*}(-\tau,-\nu)\,e^{j2\pi\nu\tau}$ at the receiver. We choose roll-off factors $\beta_{\nu}=\beta_{\tau}=0.6$ so that the effective bandwidth of the Zak-OTFS subframe is $B^{\prime}=(1+\beta_{\tau})B$ and the effective time duration is $T^{\prime}=(1+\beta_{\nu})T$. The dependence of the effective time/bandwidth and other characteristics of an OTFS modulated signal on the pulse shaping filter is discussed in detail in [24] (see equations $(23)$ and $(26)$ in [24] for the expressions of the TD and FD realizations of the Zak-OTFS pulsone).
@@
Figure 12: Uncoded $4$-QAM BER as a function of increasing $SNR$, for $\nu_{max}=1,5,9$ KHz and fixed PDR$=5$ dB.
In Fig. 12 we plot the uncoded $4$-QAM BER vs. SNR for a fixed $\nu_{max}$ and fixed PDR of $5$ dB. We observe that the conclusions made from Fig. 11 (for SNR$=25$ dB) are in fact valid over a much wider range of SNR values. Note that the BER curve floors for the single pilot case when $\nu_{max}=5,9$ KHz and it floors for the double pilot case when $\nu_{max}=9$ KHz. With four interleaved pilots, the BER curve does not floor for $\nu_{max}=1,5,9$ KHz, since the flooring would only happen when the Doppler spread $2\nu_{max}$ exceeds $304\nu_{p}=$ 30 KHz, i.e., when $\nu_{max}>15$ KHz.
Figure 13: Coded $4$-QAM block error rate (BLER) as a function of increasing $SNR$, for $\nu_{max}=1,5,9$ KHz and fixed PDR$=5$ dB.
In Fig. 13 we plot the coded block error rate (BLER) performance for Zak-OTFS modulation with the proposed interleaved pilots for acquisition of I/O relation. We use the 3GPP 5G NR Low-Density Parity Check (LDPC) codes with code-rate $1/2$ and $4-QAM$ modulation symbols. There is no coding across different Zak-OTFS frames, and therefore the code block size (number of information bits) for one, two and four interleaved pilots is $1368$, $1200$ and $868$ respectively (larger number of pilots result in a larger pilot and guard region overhead and therefore lesser number of information symbols). From the error plots, we observe error flooring for single pilot with Doppler spread more than $\nu_{p}=7.5$ KHz, i.e., $\nu_{max}>\nu_{p}/2$ ($\nu_{max}=5,9$ in the figure), and also for two interleaved pilots with Doppler spread more than $152\nu_{p}=$ 15 KHz, i.e., $\nu_{max}>7.5$ KHz ($\nu_{max}=9$ KHz in the figure). These are exactly the same scenarios where error flooring occurs in the uncoded BER plot of Fig. 12. This is because of primary reason being the same, i.e., the channel in these scenarios does not satisfy the crystallization condition w.r.t. the lattice on which the auto-ambiguity function of the interleaved pilots is supported.
@@
Figure 15: Normalized mean square error (NMSE) of the acquired channel estimates using the proposed method, as a function of increasing $\nu_{max}$ for the same setting as in Fig. 11.
In Fig. 15 we plot the normalized mean square error (NMSE) of the proposed interleaved pilots based estimation method as a function of increasing $\nu_{max}$. Let $\widehat{h_{\mbox{\scriptsize{eff}}}}[k,l]$ denote the proposed estimate of ${h_{\mbox{\scriptsize{eff}}}}[k,l]$, then the NMSE is given by ${\mathbb{E}}\left[\frac{\sum\limits_{(k,l)\in{\mathbb{S}}}|{h_{\mbox{\scriptsize{eff}}}}[k,l]-\widehat{h_{\mbox{\scriptsize{eff}}}}[k,l]|^{2}}{\sum\limits_{(k,l)\in{\mathbb{S}}}|{h_{\mbox{\scriptsize{eff}}}}[k,l]|^{2}}\right]$ where ${\mathcal{S}}$ is the support set for $h_{\mbox{\scriptsize{eff}}}[k,l]$ and the expectation is w.r.t. the random channel realizations. It is observed that with a single pilot, the estimation accuracy is poor as soon as the Doppler spread exceeds $\nu_{p}$ i.e., $\nu_{max}>\nu_{p}/2=3.75$ KHz. With two interleaved, the estimation accuracy is poor when the Doppler spread exceeds $2\nu_{p}$ and with four pilots the accuracy suffers only when the Doppler spread exceeds $4\nu_{p}$. This is due to the fact that with $Q$ interleaved pilots (regularly spaced along delay axis), the Doppler period of the auto-ambiguity function of the interleaved pilots is $Q\nu_{p}$ i.e., $Q$ times more than that for a single pilot and which is why the the cross-ambiguity between the received and the transmitted $Q$ interleaved pilot exhibits Doppler domain aliasing only when $2\nu_{max}>Q\nu_{p}$. In Figs. 12 and 13 the error rate performance floors only for those scenarios where $2\nu_{max}>Q\nu_{p}$. This confirms that the degradation in the error rate performance is primarily due to the inaccurate channel estimates (i.e., high NMSE).
@@
In Fig. 16 we plot the effective throughput as a function of increasing $\nu_{max}$ for the same simulation setting as in Fig. 11. Effective throughput is the ratio of the number of bits reliably communicated in each subframe to the available degrees of freedom ($B^{\prime}T^{\prime}$). The number of bits communicated reliably in a subframe is simply $(1-H(BER))$ times the number of information bits transmitted in each subframe. Here BER denotes the bit error rate and $H(\cdot)$ denotes the binary entropy function. We maximize effective throughput by interleaving the minimum number of pilots required to accurately estimate the effective channel. We minimize the number of pilots to avoid introducing unnecessary guard and pilot regions that would reduce effective throughput. When $2\nu_{max}<\nu_{p}$ we use a single pilot, when $\nu_{p}<2\nu_{max}<2\nu_{p}$ we use $2$ interleaved pilots, and when $2\nu_{p}<2\nu_{max}<4\nu_{p}$ we use $4$ interleaved pilots. Note that although a higher number of interleaved pilots results in stable throughput for a wider range of Doppler spreads (i.e., extension of the region of predictable operation), the throughput achieved is smaller due to a higher pilot and guard region overhead. Since $k_{max}=\lceil M\tau_{max}/\tau_{p}\rceil=2$, with single, double and four interleaved pilots, the fractional pilot overhead is $7/64$, $14/64$ and $28/64$ respectively..
Figure 17: BER vs PDR for six-path Veh-A channel used in Fig. 11 and 16, with a fixed $\nu_{max}=2.5$ KHz.
@@
The peak to average power ratio (PAPR) of the transmitted TD pilot (no data transmission) depends on the number of interleaved pilots. For the Zak-OTFS system considered here, the PAPR decreases from $19.4$ dB, to $16.4$ dB to $13.4$ dB as the number of interleaved pilots increases from $1$, to $2$, to $4$. This reduction is illustrated in Fig. 18. When the number of interleaved pilots doubles, the separation between pulses in the TD pilot pulse train halves. [^14] The TD realization of a single impulse pilot at DD location $(k_{p_{1}},0)$ is a TD pulse train with narrow TD pulses at time instances $\left(\frac{k_{p_{1}}\tau_{p}}{M}+n\tau_{p}\right)$, $n=-N/2,\cdots,0,(N/2-1)$ for even $N$ (see [19] and [1]). Therefore, the TD realization of $Q$ interleaved pilots spaced regularly apart at DD locations $(iM/Q,0)$, $i=0,1,\cdots,(Q-1)$ is a superposition of $Q$ pulse trains, each pulse train being the TD realization to one of the $Q$ pilots. The TD realization of $Q$ interleaved pilots is therefore a pulse train consisting of narrow TD pulses spaced $\tau_{p}/Q$ seconds apart. For the same total pilot energy $E_{p}$, the energy of each narrow TD pulse is therefore $E_{p}/Q$. There are twice as many pulses, and each pulse is scaled down by $\sqrt{2}$ to maintain constant average power. Hence the PAPR is halved with every doubling of the number of interleaved pilots. [^15] The PAPR of full signal (comprising both data and pilot) is smaller than that of a pilot only signal. This is because, pilot is transmitted on only a single pulsone, whereas each data symbol is transmitted on a different pulsone which reduces the peaky nature of the overall signal in TD since data pulsones located on different delay bins have peaks at different time instances. Please see Fig. $11$ in [22].
@@
In summary, with increasing number of interleaved pilots, the PAPR decreases, the guard/pilot region overhead increases and the complexity of acquiring the I/O relation also increases. As long as $\nu_{max}<Q\nu_{p}/2$, the error rate performance does not floor.
@@ VII Conclusions
Appendix A Proof of Theorem 1
Substituting the expression of $x_{p,\scriptsize{dd}}[k,l]$ from (V) into the R.H.S. of (34) we get (LABEL:axpxpkl23) (see top of next page). With $k^{\prime}=0,1\cdots,M-1$, and $l^{\prime}=0,1,\cdots,N-1$, and $i_{1}=1,2,\cdots,Q$, the Dirac-delta terms $\delta\left[k^{\prime}-(i_{1}-1)\frac{M}{Q}-n_{1}M\right]$ and $\delta\left[l^{\prime}-m_{1}N\right]$ are non-zero if and only if $l^{\prime}=n_{1}=m_{1}=0$ and $k^{\prime}=(i_{1}-1)M/Q$. With this we then get the second step in (LABEL:axpxpkl23). In the R.H.S. of the second step, the term $\delta\left[-l-m_{2}N\right]$ implies that $A_{x_{p},x_{p}}[k,l]$ is non-zero only when $l$ is an integer multiple of $N$ (i.e., $l=-m_{2}N$ for some $m_{2}\in{\mathbb{Z}}$).
Therefore, for $l=-m_{2}N$ the expression for $A_{x_{p},x_{p}}[k,l=-m_{2}N]$ is given by (LABEL:axpxpkl2) (see top of next page). From the Dirac-delta term in the R.H.S. of (LABEL:axpxpkl2) it follows that $A_{x_{p},x_{p}}[k,l=-m_{2}N]$ is non-zero only for $k=(i_{1}-i_{2})M/Q-n_{2}M$, i.e., $k$ is an integer multiple of $M/Q$. In general, $A_{x_{p},x_{p}}[k,l=-m_{2}N]$ is non-zero only for $k=pM/Q-n_{2}M$, where $p=0,1,\cdots,(Q-1)$ and $n_{2}\in{\mathbb{Z}}$. From this, it follows that for any $k\equiv pM/Q$ (modulo $M$) (i.e., $k$ is congruent to $pM/Q$ modulo $M$), we get (LABEL:axpxpkl3). From (LABEL:axpxpkl3), it is clear that $A_{x_{p},x_{p}}[k,l]$ is non-zero and equal $E_{p}$ if and only if $k$ is an integer multiple of $M/Q$ and $l$ is an integer multiple of $QN$.